OpenStudy (anonymous):

what is -x+y=5 in y=mx+b form?

OpenStudy (anonymous):

just add the x to both sides y=x+5

OpenStudy (anonymous):

do you know how that equation will be if it was perpendicular?

OpenStudy (anonymous):

Perpendicular to what?

OpenStudy (anonymous):

perpendicular to a line that has an equation of -x+y=5

OpenStudy (anonymous):

and passes through points (2,-5)

OpenStudy (anonymous):

You want the equation of a line that is perpendicular to -x+y=5 and that passes through the point (2,-5)?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Well since we put -x+y=5 into y=mx+b form y=x+5 we can see the slope is 1. a line perpendicular to that would have a slope of -1. So use that and your point to find the equation. y+5=-1(x-2) simplify that

OpenStudy (anonymous):

would it be the same if it were solving for a parallel line?

OpenStudy (anonymous):

No, parallel lines have the same slope.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

solve for y